OpenStudy (anonymous):

A 2-lb bag of brown rice costs $6. How much would you expect to pay for a 5-lb bag of brown rice, assuming there is no discount for buying a larger bag?

OpenStudy (anonymous):

The ratio is 1/3 1/3 = 5/x x = 15